### v1

# On overgroups of distinguished unipotent elements in reductive groups and finite groups of Lie type

Abstract.
Suppose $G$ is a simple algebraic group defined over an algebraically closed field of good characteristic $p$. In 2018 Korhonen showed that if $H$ is a connected reductive subgroup of $G$ which contains a distinguished unipotent element $u$ of $G$ of order $p$, then $H$ is $G$-irreducible in the sense of Serre. We present a short and uniform proof of this result using so-called good $A_{1}$ subgroups of $G$, introduced by Seitz. We also formulate a counterpart of Korhonen’s theorem for overgroups of $u$ which are finite groups of Lie type. Moreover, we generalize both results above by removing the restriction on the order of $u$ under a mild condition on $p$ depending on the rank of $G$, and we present an analogue of Korhonen’s theorem for Lie algebras.
Key words and phrases:
$G$-complete reducibility; $G$-irreducibility, distinguished unipotent elements, distinguished nilpotent elements, finite groups of Lie type, good $A_{1}$ subgroups
2010 Mathematics Subject Classification:
20G15 (14L24)
1. Introduction and main results
Throughout, $G$ is a connected reductive linear algebraic group defined over an algebraically closed field $k$ of characteristic $p\geq 0$ and $H$ is a closed subgroup of $G$.
Following Serre [30], we say that $H$ is $G$-completely reducible ($G$-cr for short) provided that whenever $H$ is contained in a parabolic subgroup $P$ of $G$, it is contained in a Levi subgroup of $P$, and that $H$ is $G$-irreducible ($G$-ir for short) provided $H$ is not contained in any proper parabolic subgroup of $G$ at all. Clearly, if $H$ is $G$-irreducible, it is trivially $G$-completely reducible, and an overgroup of a $G$-irreducible subgroup is again $G$-irreducible; for an overview of this concept see [6], [29] and [30]. Note that in case $G=\operatorname{GL}(V)$ a subgroup $H$ is $G$-completely reducible exactly when $V$ is a semisimple $H$-module and it is $G$-irreducible precisely when $V$ is an irreducible $H$-module. Recall that if $H$ is $G$-completely reducible, then the identity component $H^{\circ}$ of $H$ is reductive, [30, Prop. 4.1].
A unipotent element $u$ of $G$ is distinguished provided any torus in the centraliser $C_{G}(u)$ of $u$ in $G$ is central in $G$. Likewise, a nilpotent element $e$ of the Lie algebra $\operatorname{Lie}(G)$ of $G$ is distinguished provided any torus in the centraliser $C_{G}(e)$ of $e$ in $G$ is central in $G$, see [9, §5.9] and [15, §4.1]. For instance, regular unipotent elements in $G$ are distinguished, and so are regular nilpotent elements in $\operatorname{Lie}(G)$ [32, III 1.14] (or [9, Prop. 5.1.5]). Overgroups of regular unipotent elements have attracted much attention in the literature, e.g., see [35], [27], [37], [20], and [7].
In [16], Korhonen proves the following remarkable result in the special case when $G$ is simple.
Theorem 1.1 ([16, Thm. 6.5]).
Suppose $G$ is connected reductive and $p$ is good for $G$. Let $H$ be a reductive subgroup of $G$. Suppose $H^{\circ}$ contains a distinguished unipotent element of $G$ of order $p$. Then $H$ is $G$-irreducible.
Korhonen’s proof of Theorem 1.1 depends on checks for the various possible Dynkin types for simple $G$. E.g., for $G$ simple of exceptional type, Korhonen’s argument relies on long exhaustive case-by-case investigations from [18], where all connected reductive non-$G$-cr subgroups are classified in the exceptional type groups in good characteristic. For classical $G$, Korhonen requires an intricate classification of all $\operatorname{SL}_{2}(k)$-representations on which a non-trivial unipotent element of $\operatorname{SL}_{2}(k)$ acts with at most one Jordan block of size $p$. Our main aim is to give a short uniform proof of Theorem 1.1 in §5 without resorting to further case-by-case checks, using a landmark result by Seitz (see §4.2).
Remark 1.2.
Suppose as in Theorem 1.1, that $G$ is simple classical with natural module $V$, and $p\geq\dim V>2$. Then, thanks to [14, Prop. 3.2], $V$ is semisimple as an $H^{\circ}$-module, and by [30, (3.2.2(b))], this is equivalent to $H^{\circ}$ being $G$-cr. Then $H$ is $G$-ir, by Lemma 3.2. This gives a short uniform proof of the conclusion of Theorem 1.1 in this case, as the bound $p\geq\dim V>2$ ensures that every distinguished unipotent element (including the regular ones) is of order $p$. The conclusion can fail if the bound is not satisfied: see Theorem 1.3.
There are only a few cases when $G$ is simple, $p$ is bad for $G$, and $G$ admits a distinguished unipotent element of order $p$, by work of Proud–Saxl–Testerman [26, Lem. 4.1, Lem. 4.2] (see Lemmas 4.1 and 4.3). In this case the conclusion of Theorem 1.1 fails precisely in one instance, as observed in [16, Prop. 1.2] (Example 4.2), else it is valid (Example 4.4). Combining the cases when $p$ is bad for $G$ with Theorem 1.1, we recover Korhonen’s main theorem from [16]. Here and later on, we say that a subgroup of $G$ is of type $A_{1}$ if it is isomorphic to $\operatorname{SL}_{2}(k)$ or $\operatorname{PSL}_{2}(k)$.
Theorem 1.3 ([16, Thm. 1.3]).
Let $G$ be simple and let $H$ be a reductive subgroup of $G$. Suppose $H^{\circ}$ contains a distinguished unipotent element of $G$ of order $p$. Then $H$ is $G$-irreducible, unless $p=2$, $G$ is of type $C_{2}$, and $H$ is a type $A_{1}$ subgroup of $G$.
Our second goal is an extension of Theorem 1.1 to finite groups of Lie type in $G$. Let $\sigma:G\to G$ be a Steinberg endomorphism of $G$, so that the finite fixed point subgroup $G_{\sigma}=G(q)$ is a finite group of Lie type over the field $\mathbb{F}_{q}$ of $q$ elements. For a Steinberg endomorphism $\sigma$ of $G$ and a connected reductive $\sigma$-stable subgroup $H$ of $G$, $\sigma$ is also a Steinberg endomorphism for $H$ with finite fixed point subgroup $H_{\sigma}=H\cap G_{\sigma}$, [33, 7.1(b)]. Obviously, one cannot directly appeal to Theorem 1.1 to deduce anything about $H_{\sigma}$, because $H_{\sigma}^{\circ}$ is trivial. For the notion of a $q$-Frobenius endomorphism, see §2.4.
Theorem 1.4.
Let $H\subseteq G$ be connected reductive groups. Suppose $p$ is good for $G$. Let $\sigma\colon G\to G$ be a Steinberg endomorphism stabilizing $H$ such that $\sigma|_{H}$ is a $q$-Frobenius endomorphism of $H$. If $G$ admits components of exceptional type, then assume $q>7$. If $H_{\sigma}$ contains a distinguished unipotent element of $G$ of order $p$, then $H_{\sigma}$ is $G$-irreducible.
Combining Theorem 1.4 with the aforementioned results from [26], we are able to deduce the following analogue of Theorem 1.3 for finite subgroups of Lie type in $G$.
Theorem 1.5.
Let $G$ be simple and let $H$ be a connected reductive subgroup of $G$. Let $\sigma\colon G\to G$ be a Steinberg endomorphism stabilizing $H$ such that $\sigma|_{H}$ is a $q$-Frobenius endomorphism of $H$. If $G$ is of exceptional type, then assume $q>7$. If $H_{\sigma}$ contains a distinguished unipotent element of $G$ of order $p$, then $H_{\sigma}$ is $G$-irreducible, unless $p=2$, $G$ is of type $C_{2}$, and $H$ is a type $A_{1}$ subgroup of $G$.
In the special instance in Theorems 1.4 and 1.5 when $H_{\sigma}$ contains a regular unipotent element $u$ from $G$, the conclusion of both theorems holds without any restriction on the order of $u$ and without any restriction on $q$ (and without any exceptions of the type seen in Theorem 1.5); see [7, Thm. 1.3].
Section 2 contains background material. In Section 3 we extend Theorem 1.1 to distinguished unipotent elements of arbitrary order under an extra hypothesis on $p$, and we prove an analogue of Theorem 1.1 for Lie algebras under the same hypothesis. We recall Seitz’s notion of good $A_{1}$ subgroups in Section 4. Theorems 1.1 and 1.31.5 are proved in Section 5.
2. Preliminaries
2.1. Notation
Throughout, we work over an algebraically closed field $k$ of characteristic $p\geq 0$. All affine varieties are considered over $k$ and are identified with their $k$-points. A linear algebraic group $H$ over $k$ has identity component $H^{\circ}$; if $H=H^{\circ}$, then we say that $H$ is connected. We denote by $R_{u}(H)$ the unipotent radical of $H$; if $R_{u}(H)$ is trivial, then we say $H$ is reductive.
Throughout, $G$ denotes a connected reductive linear algebraic group over $k$. All subgroups of $G$ considered are closed. By $\mathscr{D}G$ we denote the derived subgroup of $G$, and likewise for subgroups of $G$. We denote the Lie algebra of $G$ by $\operatorname{Lie}(G)$ or by $\mathfrak{g}$.
Let $Y(G)=\operatorname{Hom}(\mathbb{G}_{m},G)$ denote the set of cocharacters of $G$. For $\mu\in Y(G)$ and $g\in G$ we define the conjugate cocharacter $g\cdot\mu\in Y(G)$ by $(g\cdot\mu)(t)=g\mu(t)g{{}^{-1}}$ for $t\in\mathbb{G}_{m}$; this gives a left action of $G$ on $Y(G)$. For $H$ a subgroup of $G$, let $Y(H):=Y(H^{\circ})=\operatorname{Hom}(\mathbb{G}_{m},H)$ denote the set of cocharacters of $H$. There is an obvious inclusion $Y(H)\subseteq Y(G)$.
Fix a Borel subgroup $B$ of $G$ containing a maximal torus $T$. Let $\Phi=\Phi(G,T)$ be the root system of $G$ with respect to $T$, let $\Phi^{+}=\Phi(B,T)$ be the set of positive roots of $G$, and let $\Sigma=\Sigma(G,T)$ be the set of simple roots of $\Phi^{+}$. For each $\alpha\in\Phi$ we have a root subgroup $U_{\alpha}$ of $G$. For $\alpha$ in $\Phi$, let $x_{\alpha}:\mathbb{G}_{a}\to U_{\alpha}$ be a parametrization of the root subgroup $U_{\alpha}$ of $G$.
2.2. Good primes
A prime $p$ is said to be good for $G$ if it does not divide any coefficient of any positive root when expressed as a linear combination of simple ones. Else $p$ is called bad for $G$, [32, §4]. Explicitly, if $G$ is simple, $p$ is good for $G$ provided $p>2$ in case $G$ is of Dynkin type $B_{n}$, $C_{n}$, or $D_{n}$; $p>3$ in case $G$ is of Dynkin type $E_{6}$, $E_{7}$, $F_{4}$ or $G_{2}$ and $p>5$ in case $G$ is of type $E_{8}$. We observe that if $L$ is a Levi subgroup of $G$ and $p$ is good for $G$, then it is also good for $L$.
2.3. Springer isomorphisms
If $p$ is good for $G$ and $\mathscr{D}G$ is simply connected, then there exists a $G$-equivariant homeomorphism $\phi:\mathscr{U}\to\mathscr{N}$ between the unipotent variety $\mathscr{U}$ of $G$ and the nilpotent cone $\mathscr{N}$ of $\operatorname{Lie}(G)$. Such a map is called a Springer isomorphism, see [32, III, 3.12] and [4, Cor. 9.3.4]. By means of such a map, all the results below formulated for unipotent elements in $G$ stem from analogues for nilpotent elements in $\operatorname{Lie}(G)$. In what follows we fix such a Springer isomorphism once and for all.
2.4. Steinberg endomorphisms of $G$
[…]
If $\sigma_{q}:G\to G$ is a standard $q$-power Frobenius endomorphism of $G$, then there exists a $\sigma_{q}$-stable maximal torus $T$ and Borel subgroup $B\supseteq T$, and with respect to a chosen parametrisation of the root groups as above, we have $\sigma_{q}(x_{\alpha}(t))=x_{\alpha}(t^{q})$ for each $\alpha\in\Phi$ and $t\in\mathbb{G}_{a}$, see [12, Thm. 1.15.4(a)]. Following [26], we call a generalized Frobenius endomorphism $\sigma$ a $q$-Frobenius endomorphism provided $\sigma=\tau\sigma_{q}$, where $\tau$ is an algebraic automorphism of $G$ of finite order, $\sigma_{q}$ is a standard $q$-power Frobenius endomorphism of $G$, and $\sigma_{q}$ and $\tau$ commute. When $p$ is bad for $G$, a $q$-Frobenius endomorphism does not involve a twisted Steinberg endomorphism, see [26, §3]. In particular, if $G$ is simple and $p$ is good for $G$, then any Steinberg endomorphism of $G$ is a $q$-Frobenius endomorphism, see [33, §11]. If $G$ is not simple and $p$ is bad for $G$, then a generalized Frobenius map may fail to factor into a field and algebraic automorphism of $G$, e.g. see [13, Ex. 1.3].
2.5. Bala–Carter Theory
We recall some relevant results and concepts from Bala–Carter Theory. A parabolic subgroup $P$ of $G$ admits a dense open orbit on its unipotent radical $R_{u}(P)$, the so-called Richardson orbit; see [9, Thm. 5.2.1]. A parabolic subgroup $P$ of $G$ is called distinguished provided $\dim(\mathscr{D}P/R_{u}(P))=\dim(R_{u}(P)/\mathscr{D}R_{u}(P))$, see [9, §5.8]. For $G$ simple, the distinguished parabolic subgroups of $G$ (up to $G$-conjugacy) were worked out in [2] and [3]; see [9, pp. 174–177].
The following is the celebrated Bala–Carter theorem, see [9, Thm. 5.9.5, Thm. 5.9.6], which is valid in good characteristic, thanks to work of Pommerening [23], [24]. For the Lie algebra versions see also [15, Prop. 4.7, Thm. 4.13].
Theorem 2.1.
Suppose $p$ is good for $G$.
  • (i)
    There is a bijective map between the $G$-conjugacy classes of distinguished unipotent elements of $G$ and conjugacy classes of distinguished parabolic subgroups of $G$. The unipotent class corresponding to a given parabolic subgroup $P$ contains the dense $P$-orbit on $R_{u}(P)$.
  • (ii)
    There is a bijective map between the $G$-conjugacy classes of unipotent elements of $G$ and conjugacy classes of pairs $(L,P)$, where $L$ is a Levi subgroup of $G$ and $P$ is a distinguished parabolic subgroup of $\mathscr{D}L$. The unipotent class corresponding to the pair $(L,P)$ contains the dense $P$-orbit on $R_{u}(P)$.
Remark 2.2.
(i). Let $u\in G$ be unipotent. Let $S$ be a maximal torus of $C_{G}(u)$. Then $u$ is distinguished in the Levi subgroup $C_{G}(S)$ of $G$, for $S$ is the unique maximal torus of $C_{C_{G}(S)}(u)$. Conversely, if $L$ is a Levi subgroup of $G$ with $u$ distinguished in $L$, then the connected center of $L$ is a maximal torus of $C_{G}(u)^{\circ}$, see [15, Rem. 4.7].
(ii). Let $\sigma:G\to G$ be a Steinberg endomorphism of $G$ and let $u\in G_{\sigma}$ be unipotent. Then $C_{G}(u)^{\circ}$ is $\sigma$-stable. The set of all maximal tori of $C_{G}(u)^{\circ}$ is $\sigma$-stable and $C_{G}(u)^{\circ}$ is transitive on that set, [31, Thm. 6.4.1]. Thus the Lang–Steinberg Theorem, see [32, I 2.7], provides a $\sigma$-stable maximal torus, say $S$, of $C_{G}(u)^{\circ}$. Then, by part (i), $L=C_{G}(S)$ is a $\sigma$-stable Levi subgroup of $G$ and $u$ is distinguished in $L$.
2.6. Cocharacters and parabolic subgroups of $G$
Let $\lambda\in Y(G)$. Without loss, we may assume that $\lambda\in Y(T)$. Recall that $\lambda$ affords a $\mathbb{Z}$-grading on $\operatorname{Lie}(G)=\mathfrak{g}=\bigoplus_{j\in\mathbb{Z}}\mathfrak{g}(j,\lambda)$, where $\mathfrak{g}(j,\lambda):=\{e\in\mathfrak{g}\mid\operatorname{Ad}(\lambda(t))e=t^{j}e\text{ for every }t\in\mathbb{G}_{m}\}$ is the $j$-weight space of $\operatorname{Ad}(\lambda(\mathbb{G}_{m}))$ on $\mathfrak{g}$, see [9, §5.5] or [15, §5.1]. Let $\mathfrak{p}_{\lambda}:=\bigoplus_{j\geq 0}\mathfrak{g}(j,\lambda)$. Then there is a unique parabolic subgroup $P_{\lambda}$ with $\operatorname{Lie}P_{\lambda}=\mathfrak{p}_{\lambda}$ and $C_{G}(\lambda):=C_{G}(\lambda(\mathbb{G}_{m}))$ is a Levi subgroup of $P_{\lambda}$. Specifically, we have $U_{\alpha}\subseteq P_{\lambda}$ if and only if $\langle\lambda,\alpha\rangle\geq 0$, where $\langle\,,\,\rangle:Y(T)\times X(T)\to\mathbb{Z}$ is the usual pairing between cocharacters and characters $X(T)=\operatorname{Hom}(T,\mathbb{G}_{m})$ of $T$. We have $U_{\alpha}\subseteq C_{G}(\lambda)$ if and only if $\langle\lambda,\alpha\rangle=0$, and $R_{u}(P_{\lambda})$ is generated by the $U_{\alpha}$ with $\langle\lambda,\alpha\rangle>0$; see the proof of [31, Prop. 8.4.5].
Set $J:=\{\alpha\in\Sigma\mid\langle\alpha,\lambda\rangle=0\}$. Then $P_{\lambda}=P_{J}=\big{\langle}T,U_{\alpha}\mid\langle\alpha,\lambda\rangle\geq 0\big{\rangle}$ is the standard parabolic subgroup of $G$ associated with $J\subseteq\Sigma$.
Let $\rho=\sum_{\alpha\in\Sigma}c_{\alpha\rho}\alpha$ be the highest root in $\Phi^{+}$. Define $\operatorname{ht}_{J}(\rho):=\sum_{\alpha\in\Sigma\setminus J}c_{\alpha\rho}$. In view of Theorem 2.1, the following gives the order of a distinguished unipotent element in good characteristic.
Lemma 2.3 ([36, Order Formula 0.4]).
Let $p$ be good for $G$. Let $P=P_{J}$ be a distinguished parabolic subgroup of $G$ and let $u$ be in the Richardson orbit of $P$ on $R_{u}(P)$. Then the order of $u$ is $\min\{p^{a}\mid p^{a}>\operatorname{ht}_{J}(\rho)\}$.
2.7. Cocharacters associated to nilpotent and unipotent elements
The Jacobson–Morozov Theorem allows one to associate an $\mathfrak{sl}(2)$-triple to any given non-zero nilpotent element in $\mathscr{N}$ in characteristic zero or large positive characteristic. This is an indispensable tool in the Dynkin–Kostant classification of the nilpotent orbits in characteristic zero as well as in the Bala–Carter classification of unipotent conjugacy classes of $G$ in large prime characteristic, see [9, §5.9]. In good characteristic there is a replacement for $\mathfrak{sl}(2)$-triples, so-called associated cocharacters; see Definition 2.4 below. These cocharacters are important tools in the classification theory of unipotent and nilpotent classes of reductive algebraic groups in good characteristic, see for instance [15, §5] and [25]. We recall the relevant concept of cocharacters associated to a nilpotent element following [15, §5.3].
Definition 2.4.
A cocharacter $\lambda\in Y(G)$ of $G$ is associated to $e\in\mathscr{N}$ provided $e\in\mathfrak{g}(2,\lambda)$ and there exists a Levi subgroup $L$ of $G$ such that $e$ is distinguished nilpotent in $\operatorname{Lie}L$ and $\lambda(\mathbb{G}_{m})\leq\mathscr{D}L$.
For the remainder of this section suppose that $\mathscr{D}G$ is simply connected. A cocharacter $\lambda\in Y(G)$ of $G$ is associated to $u\in\mathscr{U}$ provided it is associated to $\phi(u)$, where $\phi:\mathscr{U}\to\mathscr{N}$ is a fixed Springer isomorphism as in §2.3; see [22, Rem. 23]. Following [11, Def. 2.13], we write
$\Omega^{a}_{G}(u):=\{\lambda\in Y(G)\mid\lambda\text{ is associated to }u\}$
for the set of cocharacters of $G$ associated to $u$. This notation stems from the fact that associated cocharacters are destabilising cocharacters of $G$ for $u$ in the sense of Kempf–Rousseau theory, see [25] and [22]. Likewise, for $M$ a connected reductive subgroup of $G$ containing $u$, we write $\Omega^{a}_{M}(u)$ for the set of cocharacters of $M$ that are associated to $u$.
Remark 2.5.
Let $u\in G$ be unipotent, $\lambda\in\Omega^{a}_{G}(u)$, and $g\in C_{G}(u)$. Then $g\cdot\lambda$ is also associated to $u$, see [15, §5.3]. Proposition 2.6(ii) gives a converse to this property.
We require some basic facts about cocharacters associated to unipotent elements. The following results are [15, Lem. 5.3; Prop. 5.9] for nilpotent elements, see also [25, Thm. 2.3, Prop. 2.5].
Proposition 2.6.
Suppose $p$ is good for $G$. Let $u\in G$ be unipotent.
  • (i)
    $\Omega^{a}_{G}(u)\neq\varnothing$, i.e., cocharacters of $G$ associated to $u$ exist.
  • (ii)
    $C_{G}(u)^{\circ}$ acts transitively on $\Omega^{a}_{G}(u)$.
  • (iii)
    Let $\lambda\in\Omega^{a}_{G}(u)$ and let $P_{\lambda}$ be the parabolic subgroup of $G$ defined by $\lambda$ as in §2.6. Then $P_{\lambda}$ only depends on $u$ and not on the choice of $\lambda$.
  • (iv)
    Let $\lambda\in\Omega^{a}_{G}(u)$ and let $P(u):=P_{\lambda}$ be as in (iii). Then $C_{G}(u)\subseteq P(u)$.
If $u$ is distinguished in $G$, then the parabolic subgroup $P(u)$ of $G$ from Proposition 2.6(iii) is a distinguished parabolic subgroup of $G$ and $u$ belongs to the Richardson orbit of $P(u)$ on its unipotent radical, see Theorem 2.1(i); see also [22, Prop. 22].
Let $u$ be unipotent in $G$ and let $\lambda\in\Omega^{a}_{G}(u)$. Let $P=P(u)$ be the canonical parabolic subgroup defined by $u$ from Proposition 2.6. Then $P=C_{G}(\lambda)R_{u}(P)$ is a Levi decomposition of $P$. Following [15, §5.10] and [25, §2.4], we define the subgroup
(2.7) $C_{G}(u,\lambda):=C_{G}(u)\cap C_{G}(\lambda)$
of $C_{G}(u)$. In view of [25, Prop. 2.5], our next result is [25, Thm. 2.3(iii)], see also [15, Prop. 5.10, Prop. 5.11].
Proposition 2.8.
Suppose $p$ is good for $G$. Let $u\in G$ be unipotent and let $\lambda\in\Omega^{a}_{G}(u)$. Then $C_{G}(u)$ is the semidirect product of $C_{G}(u,\lambda)$ and $R_{u}(C_{G}(u))$, and $C_{G}(u,\lambda)$ is reductive.
We note that cocharacters of $G$ associated to a unipotent element $u$ of $G$ are compatible with certain group-theoretic operations; see [15, §5.6].
Remark 2.9.
Suppose $p>0$ and $u\in G$ is unipotent of order $p$ contained in a subgroup $A$ of $G$ of type $A_{1}$. Such a subgroup $A$ always exists when $p$ is good, and when $p$ is bad there is essentially only one exception, due to Testerman [36] and Proud–Saxl–Testerman [26] — see Theorems 4.5 and 4.6 below. Then, since $p$ is good for $A$, by Proposition 2.6(i) there exists a cocharacter $\lambda\in\Omega^{a}_{A}(u)$. Note that $\lambda(\mathbb{G}_{m})$ is a maximal torus in $A$.
It follows from the work of Pommerening [23], [24] that the description of the unipotent classes in characteristic 00 is identical to the one for $G$ when $p$ is good for $G$. In both instances these are described by so-called weighted Dynkin diagrams. As a result, an associated cocharacter to a unipotent element acts with the very same weights on the Lie algebra of $G$ as its counterpart does in characteristic 00. This fact is used in the proof of the following result by Lawther [17, Thm. 1]; see also the proof of [28, Prop. 4.2].
Lemma 2.10.
Let $G$ be simple and suppose $u\in G$ is unipotent. Suppose $p$ is good for $G$. Let $\lambda\in\Omega^{a}_{G}(u)$. Denote by $\omega_{G}$ the highest weight of $\lambda(\mathbb{G}_{m})$ on $\operatorname{Lie}(G)$. Then $u$ has order $p$ if and only if $\omega_{G}\leq 2p-2$.
The concept of associated cocharacter is not only a convenient replacement for $\mathfrak{sl}(2)$-triples from the Jacobson–Morozov Theory, it is a very powerful tool in the classification theory of unipotent conjugacy classes and nilpotent orbits. Specifically, in [25] Premet showcases a conceptual and uniform proof of Pommerening’s extension of the Bala–Carter Theorem 2.1 to good characteristic. His proof uses the fact that associated cocharacters are optimal in the geometric invariant theory sense of Kempf–Rousseau–Hesselink. In §5, we demonstrate the utility of this concept further with short uniform proofs of parts of Seitz’s Theorem 4.10 on good $A_{1}$ subgroups.
2.8. Cocharacters associated to distinguished elements
The linchpin of our proofs of Theorems 1.1 and 1.4 is the following fact.
Lemma 2.11 ([11, Lem. 3.1]).
Suppose $p$ is good for $G$. Let $M\subset G$ be a connected reductive subgroup of $G$. Let $u\in M$ be a distinguished unipotent element of $G$. Then $\Omega^{a}_{M}(u)=\Omega^{a}_{G}(u)\cap Y(M)$.
The assertion of the lemma fails in general if $u$ is not distinguished in $G$, even when $p$ is good for both $M$ and $G$, e.g., see [15, Rem. 5.12]. However, we do have the following result for all unipotent elements in good characteristic.
Lemma 2.12 ([11, Cor. 3.22]).
Suppose $p$ is good for $G$. Let $L\subset G$ be a Levi subgroup of $G$. Let $u\in L$ be unipotent. Then $\Omega^{a}_{L}(u)=\Omega^{a}_{G}(u)\cap Y(L)$.
3. Generalizations of Theorems 1.1 and 1.4
The aim of this section is to generalize Theorem 1.1 to the situation where there is no restriction on the order of the unipotent element at hand, see Theorem 3.1. For a unipotent element $u$ in $G$ to be distinguished is a mere condition on the structure of the centralizer $C_{G}(u)$ of $u$ in $G$. The extra condition for $u$ to have order $p$ is thus somewhat artificial. The restriction in Theorem 1.1 is due to the methods used in [16] and in our proofs in §5, which require the unipotent element to lie in a subgroup of type $A_{1}$; such an element must obviously have order $p$.
Along the way we prove an analogue of Theorem 1.1 for Lie algebras, i.e., under the hypothesis that $\operatorname{Lie}(H)$ contains a distinguished nilpotent element of $\operatorname{Lie}(G)$. In order to state this theorem, we need to introduce an invariant $a(G)$ of $G$ from [30, §5.2]: for $G$ simple, set $a(G)=\operatorname{rk}(G)+1$, where $\operatorname{rk}(G)$ is the rank of $G$. For reductive $G$, let $a(G)=\max\{1,a(G_{1}),\ldots,a(G_{r})\}$, where $G_{1},\ldots,G_{r}$ are the simple components of $G$.
Theorem 3.1.
Suppose $p\geq a(G)$. Let $H$ be a reductive subgroup of $G$. Suppose $H^{\circ}$ contains a distinguished unipotent element of $G$ or $\operatorname{Lie}(H)$ contains a distinguished nilpotent element of $\operatorname{Lie}(G)$. Then $H$ is $G$-irreducible.
The following analogue of [7, Cor. 4.6] shows that in order to derive the $G$-irreducibility of $H$ in Theorem 3.1, it suffices to show that $H$ is $G$-cr, see also [16, Lem. 6.1]. This also applies to Theorem 1.1 and Theorem 1.4.
Lemma 3.2.
Let $H$ be a $G$-completely reducible subgroup of $G$. Suppose that $H$ contains a distinguished unipotent element $u$ of $G$ or $\operatorname{Lie}(H)$ contains a distinguished nilpotent element $e$ of $\operatorname{Lie}(G)$. Then $H$ is $G$-irreducible.
Proof.
Suppose $H$ is contained in a parabolic subgroup $P$ of $G$. Then, by hypothesis, $H$ is contained in a Levi subgroup $L$ of $P$. As the latter is the centraliser of a torus $S$ in $G$, $S$ centralises $u$ (resp., $e$) and so $S$ is central in $G$. Hence $L=G$, which implies $P=G$. ∎
Along with Lemma 3.2, the following theorem of Serre immediately yields Theorem 3.1.
Theorem 3.3 ([30, Thm. 4.4]).
Suppose $p\geq a(G)$ and $(H:H^{\circ})$ is prime to $p$. Then $H^{\circ}$ is reductive if and only if $H$ is $G$-completely reducible.
Proof of Theorem 3.1.
Since $p\geq a(G)$, Theorem 3.3 applied to $H^{\circ}$ shows the latter is $G$-cr. Thus $H^{\circ}$ is $G$-ir by Lemma 3.2, and so is $H$. ∎
Remarks 3.4.
(i). The characteristic restriction in Theorem 3.1 (and Theorem 3.3) is needed; see Theorem 1.3.
(ii). The condition in Theorem 3.1 requiring $H^{\circ}$ to contain a distinguished unipotent element of $G$ (as opposed to $H$) is also necessary, as for instance the finite unipotent subgroup of $G$ generated by a given distinguished unipotent element of $G$ is not $G$-cr [30, Prop. 4.1].
(iii). Under the given hypotheses, Theorem 3.1 applies to an arbitrary distinguished unipotent element of $G$, irrespective of its order. For Theorem 1.1 to achieve the same uniform result, $p$ has to be sufficiently large to guarantee that the chosen element has order $p$. For $G$ simple classical with natural module $V$, this requires the bound $p\geq\dim V$; see Remark 1.2. For $G$ simple of exceptional type, this requires the following bounds: $p>11$ for $E_{6}$, $p>17$ for $E_{7}$, $p>29$ for $E_{8}$, $p>11$ for $F_{4}$, and $p>5$ for $G_{2}$; see [36, Prop. 2.2]. So in many cases the bound $p\geq a(G)$ from Theorem 3.1 is better.
(iv). For an instance when $p$ is bad for $G$ so that Theorem 1.1 does not apply, but Theorem 3.1 does, see Example 4.4.
(v). Theorem 3.1 generalizes [7, Thm. 3.2] which consists of the analogue in the special instance when the distinguished element is regular in $G$ (or $\operatorname{Lie}(G)$). Note that in this case no restriction on $p$ is needed, see [37, Thm. 1.2], [20, Thm. 1], [7, Thm. 3.2].
(vi). In characteristic 00, a subgroup $H$ of $G$ is $G$-cr if and only if it is reductive, [30, Prop. 4.1]. So in that case the conclusion of Theorem 3.1 follows directly from Lemma 3.2.
Once again, in the presence of a Steinberg endomorphism $\sigma$ of $G$, one cannot appeal to Theorem 3.1 directly to deduce anything about $H_{\sigma}$, because $H_{\sigma}^{\circ}$ is trivial. In Corollary 3.6 we present an analogue of Theorem 3.1 for the finite groups of Lie type $H_{\sigma}$ under an additional condition stemming from [5].
Note that for $S$ a torus in $G$, we have $C_{G}(S)=C_{G}(s)$ for some $s\in S$, see [8, III Prop. 8.18].
Proposition 3.5 ([5, Prop. 3.2]).
Let $H\subseteq G$ be connected reductive groups. Let $\sigma\colon G\to G$ be a Steinberg endomorphism that stabilises $H$ and a maximal torus $T$ of $H$. Suppose
  • (i)
    $C_{G}(T)=C_{G}(t)$, for some $t\in T_{\sigma}$, and
  • (ii)
    $H_{\sigma}$ meets every $T$-root subgroup of $H$ non-trivially.
Then $H_{\sigma}$ and $H$ belong to the same parabolic and the same Levi subgroups of $G$. In particular, $H$ is $G$-completely reducible if and only if $H_{\sigma}$ is $G$-completely reducible; similarly, $H$ is $G$-irreducible if and only if $H_{\sigma}$ is $G$-irreducible.
Without condition (i), the proposition is false in general, see [5, Ex. 3.2]. The following is an immediate consequence of Theorem 3.1 and Proposition 3.5.
Corollary 3.6.
Suppose $G,H$ and $\sigma$ satisfy the hypotheses of Proposition 3.5. Suppose in addition that $p\geq a(G)$. If $H_{\sigma}$ contains a distinguished unipotent element of $G$, then $H_{\sigma}$ is $G$-irreducible.
Corollary 3.6 generalizes [7, Thm. 1.3] which consists of the analogue in the special instance when the distinguished element is regular in $G$. Note in this case no restriction on $p$ is needed.
The following example shows that the conditions in Corollary 3.6 hold generically.
Example 3.7.
Let $\sigma_{q}\colon\operatorname{GL}(V)\to\operatorname{GL}(V)$ be a standard Frobenius endomorphism which stabilises a connected reductive subgroup $H$ of $\operatorname{GL}(V)$ and a maximal torus $T$ of $H$. Pick $l\in\mathbb{N}$ such that firstly all the different $T$-weights of $V$ are still distinct when restricted to $T_{\sigma_{q}^{l}}$ and secondly there is a $t\in T_{\sigma_{q}^{l}}$, such that $C_{\operatorname{GL}(V)}(T)=C_{\operatorname{GL}(V)}(t)$. Then for every $n\geq l$, both conditions in Corollary 3.6 are satisfied for $\sigma=\sigma_{q}^{n}$. Thus there are only finitely many powers of $\sigma_{q}$ for which part (i) can fail. The argument here readily generalises to a Steinberg endomorphism of a connected reductive $G$ which induces a generalised Frobenius morphism on $H$.
4. Overgroups of type $A_{1}$
4.1. Overgroups of type $A_{1}$
We begin with the distinguished case. There are only a few instances when $G$ is simple, $p$ is bad for $G$, and $G$ admits a distinguished unipotent element of order $p$. We recall the relevant results concerning the existence of $A_{1}$ overgroups of such elements from [26].
Lemma 4.1 ([26, Lem. 4.1]).
Let $G$ be simple classical of type $B_{l},C_{l}$, or $D_{l}$ and suppose $p=2$. Then $G$ admits a distinguished involution $u$ if and only if $G$ is of type $C_{2}$ and $u$ belongs to the subregular class $\mathscr{C}$ of $G$. If $\sigma$ is $\operatorname{id}_{G}$ or a $q$-Frobenius endomorphism of $G$ and $u\in\mathscr{C}\cap G_{\sigma}$, then there exists a $\sigma$-stable subgroup $A$ of $G$ of type $A_{1}$ containing $u$.
Example 4.2.
Let $G$ be simple of type $C_{2}$, $p=2$, and suppose $u$ is a distinguished unipotent of order $2$ in $G$. Let $\sigma$ be $\operatorname{id}_{G}$ or a $q$-Frobenius endomorphism of $G$. Then Lemma 4.1 provides a $\sigma$-stable subgroup $A$ of type $A_{1}$ containing $u$. Thanks to [16, Prop. 1.2], there are such subgroups $A$ which are not $G$-ir. In fact, according to loc. cit., there are two $G$-conjugacy classes of such $A_{1}$ subgroups in $G$; see also Example 4.17 below. Since $A$ is contained in a proper parabolic subgroup of $G$, so is $A_{\sigma}$. So the latter is also not $G$-ir.
Lemma 4.3 ([26, Lem. 3.3, Lem. 4.2]).
Let $G$ be simple of exceptional type and suppose $p$ is bad for $G$. Then $G$ admits a distinguished unipotent element $u$ of order $p$ if and only if $G$ is of type $G_{2}$, $p=3$, and either $u$ belongs to the subregular class $G_{2}(a_{1})$ [^1] Throughout, we use the Bala–Carter notation for distinguished classes in the exceptional groups, see [9, §5.9]. or to the class $A_{1}^{(3)}$ of $G$. Moreover, if $\sigma$ is $\operatorname{id}_{G}$ or a $q$-Frobenius endomorphism of $G$ and $u\in G_{2}(a_{1})\cap G_{\sigma}$, then there exists a $\sigma$-stable subgroup $A$ of $G$ of type $A_{1}$ containing $u$. In case $u\in A_{1}^{(3)}$, there is no overgroup of $u$ in $G$ of type $A_{1}$.
Example 4.4.
Let $G$ be simple of type $G_{2}$ and $p=3$. Let $H$ be a reductive subgroup of $G$ containing a distinguished unipotent element $u$ from $G$. Then, as $p=3=a(G_{2})$, it follows from Theorem 3.1 that $H^{\circ}$ is $G$-ir, and so is $H$. This applies in particular to the subgroup $A$ of $G$ of type $A_{1}$ containing $u$ when $u\in G_{2}(a_{1})$. Since $3$ is not a good prime for $G$, Theorem 1.1 does not apply in this case. See also [34, Cor. 2].
In case of the presence of a $q$-Frobenius endomorphism of $G$ stabilizing $H$, we show in our proof of Theorem 1.5 that $H_{\sigma}$ is also $G$-ir.
The existence of $A_{1}$ overgroups for unipotent elements of order $p$ is guaranteed by the following fundamental results of Testerman [36, Thm. 0.1] if $p$ is good for $G$ and else by Proud–Saxl–Testerman [26].
Theorem 4.5 ([36, Thm. 0.1, Thm. 0.2]).
Let $G$ be a semisimple group and suppose $p$ is good for $G$. Let $\sigma$ be $\operatorname{id}_{G}$ or a Steinberg endomorphism of $G$. Let $u\in G_{\sigma}$ be unipotent of order $p$. Then there exists a $\sigma$-stable subgroup of $G$ of type $A_{1}$ containing $u$.
Testerman’s proof of Theorem 4.5 is based on case-by-case checks and depends in part on computer calculations involving explicit unipotent class representatives. For a uniform proof of the theorem, we refer the reader to McNinch [21].
Theorem 4.6 ([26, Thm. 5.1]).
Let $G$ be semisimple and suppose $p$ is bad for $G$. Let $\sigma$ be $\operatorname{id}_{G}$ or a $q$-Frobenius endomorphism of $G$. Let $u\in G_{\sigma}$ be unipotent of order $p$. If $p=3$, and $G$ has a simple component of type $G_{2}$, assume that the projection of $u$ into this component does not lie in the class $A_{1}^{(3)}$. Then there exists a $\sigma$-stable subgroup of $G$ of type $A_{1}$ containing $u$.
Corollary 4.7.
Let $G$ be simple of type $G_{2}$, $p=3$ and let $\sigma$ be $\operatorname{id}_{G}$ or a $q$-Frobenius endomorphism of $G$. Let $u\in A_{1}^{(3)}\cap G_{\sigma}$. Then there is no proper semisimple subgroup $H$ of $G$ containing $u$. In particular, any such $u$ is semiregular, that is, $C_{G}(u)$ does not contain a non-central semisimple element of $G$.
Proof.
By way of contradiction, suppose $H$ is a proper semisimple subgroup of $G$ containing $u$. Since $p=3$ is good for $H$ (e.g., see [34, Cor. 3]), there is a $\sigma$-stable $A_{1}$ subgroup $A$ in $H$ containing $u$, by Theorem 4.5. It follows from Lemma 4.3 that $u\in G_{2}(a_{1})$ which contradicts the hypothesis that $u\in A_{1}^{(3)}$. ∎
With Theorem 4.6 in hand, we need to deal next with the case when $p=3$ is good for $G$ and $H$ admits a simple component of type $G_{2}$.
Lemma 4.8.
Suppose $G$ is simple and $p=3$ is good for $G$. Let $H$ be a connected reductive subgroup of $G$. Let $u\in H$ be a unipotent element of order $3$. Then $H$ does not admit a simple component of type $G_{2}$.
Proof.
([16, p. 387]) Since $p$ is good for $G$, $G$ is simple of classical type. Let $V$ be the natural module for $G$. Since $u$ has order $3$, the largest Jordan block size of $u$ on $V$ is at most $3$. Since $u$ is distinguished in $G$, the Jordan block sizes of $u$ are distinct and of the same parity. Hence $\dim V\leq 4$. Since a non-trivial representation of a simple algebraic group of type $G_{2}$ has dimension at least $5$, $H$ does not have a simple component of type $G_{2}$. ∎
In summary, we see that if $u\in G$ has order $p$ then $u$ is contained in an $A_{1}$ subgroup of $G$ unless $p=3$ and $G$ has a simple $G_{2}$ factor such that the projection of $u$ onto this factor lies in the class $A_{1}^{(3)}$.
4.2. Good $A_{1}$ overgroups
In his seminal work [28], Seitz defines an important class of $A_{1}$ overgroups of a unipotent element of order $p$. He establishes the existence and fundamental properties of these overgroups provided $p$ is good for $G$.
Definition 4.9.
Following [28], we call a subgroup $A$ of $G$ of type $A_{1}$ a good $A_{1}$ subgroup of $G$ if the weights of a maximal torus of $A$ on $\operatorname{Lie}(G)$ are at most $22p-$ 2. Else we call $A$ a bad $A_{1}$ subgroup of $G$. This is of course independent of the choice of a maximal torus of $A$. For $u$ a unipotent element of $G$ of order $p$, we define
$\mathscr{A}(u):=\mathscr{A}_{G}(u):=\{A\subseteq G\mid\text{$A$ is a good $A_{1}$ subgroup of $G$ containing $u$}\}$
and analogously, for a connected reductive subgroup $M$ of $G$ we write $\mathscr{A}_{M}(u)$ for the set of all good $A_{1}$ subgroups of $M$ containing $u$.
If $A\subseteq H\subseteq G$ are connected reductive groups such that $A$ is a good $A_{1}$ in $G$, then $A$ is obviously also a good $A_{1}$ in $H$.
We note that the property of being a good $A_{1}$ subgroup is unaffected by central isogenies (in both directions). This is clear, because Definition 4.9 is concerned with the nontrivial weights of a torus of the $A_{1}$ subgroup; the associated weight spaces consist of nilpotent elements of $\operatorname{Lie}(G)$, and central isogenies give rise to equivariant bijections on the sets of nilpotent elements; cf. [25, §2.3].
We record parts of the main theorems from [28] for our purposes, using the notation above.
Theorem 4.10 ([28, Thm. 1.1, Thm. 1.2]).
Let $G$ be simple. Suppose $p$ is good for $G$ and $u$ is a unipotent element of $G$ of order $p$. Then the following hold:
  • (i)
    $\mathscr{A}(u)\neq\varnothing$.
  • (ii)
    $R_{u}(C_{G}(u))$ acts transitively on $\mathscr{A}(u)$.
  • (iii)
    Let $A\in\mathscr{A}(u)$. Then $C_{G}(u)$ is the semidirect product of $C_{G}(A)$ and $R_{u}(C_{G}(u))$ and $C_{G}(A)$ is reductive.
  • (iv)
    Let $A\in\mathscr{A}(u)$. Then $A$ is $G$-completely reducible.
The following example is a special instance of Theorem 4.10.
Example 4.11.
Let $V$ be an $\operatorname{SL}_{2}(k)$-module such that weights of a maximal torus $T$ of $\operatorname{SL}_{2}(k)$ on $V$ are strictly less than $p$. Then the weights of $T$ in the induced action on $\operatorname{Lie}(\operatorname{GL}(V))\cong V\otimes V^{*}$ are at most $22p-$ 2. Thus the induced subgroup $A$ in $\operatorname{GL}(V)$ is a good $A_{1}$. In this situation the highest weights of $T$ on each composition factor of $V$ are restricted, so $V$ is a semisimple $\operatorname{SL}_{2}(k)$-module; see [1, Cor. 3.9].
In the next theorem we recall parts of the analogue of Theorem 4.10 for finite overgroups of type $A_{1}$.
Theorem 4.12 ([28, Thm. 1.4]).
Let $G$ be simple. Suppose $p$ is good for $G$. Let $\sigma:G\to G$ be a Steinberg endomorphism of $G$. Suppose $u\in G_{\sigma}$ is unipotent of order $p$.
  • (i)
    $\mathscr{A}(u)_{\sigma}\neq\varnothing$.
  • (ii)
    $R_{u}(C_{G}(u))_{\sigma}$ acts transitively on $\mathscr{A}(u)_{\sigma}$.
  • (iii)
    Let $A\in\mathscr{A}(u)_{\sigma}$. Suppose that $q>7$ if $G$ is of exceptional type. Then $A_{\sigma}$ is $\sigma$-completely reducible.
Remark 4.13.
Parts (i) and (ii) of Theorem 4.12 follow from parts (i) and (ii) of Theorem 4.10 and the Lang–Steinberg Theorem, see [28, Prop. 9.1].
Remark 4.14.
(i). Concerning the terminology in Theorem 4.12(iii), following [13], a subgroup $H$ of $G$ is said to be $\sigma$-completely reducible, provided that whenever $H$ lies in a $\sigma$-stable parabolic subgroup $P$ of $G$, it lies in a $\sigma$-stable Levi subgroup of $P$. This notion is motivated by certain rationality questions concerning $G$-complete reducibility; see [13] for details. For a $\sigma$-stable subgroup $H$ of $G$, this property is equivalent to $H$ being $G$-cr, thanks to [13, Thm. 1.4]. Thus in turn, by the latter, Theorem 4.12(iii) follows from Theorem 4.10(iv).
(ii). Apart from the special conjugacy class of good $A_{1}$ subgroups in $G$ asserted in Theorem 4.10, there might be a plethora of conjugacy classes of bad $A_{1}$ subgroups in $G$ even when $p$ is good for $G$. Just take a non-semisimple representation $\rho:\operatorname{SL}_{2}(k)\to\operatorname{SL}(V)=G$ in characteristic $p>0$. Then the $A_{1}$ subgroup $\rho(\operatorname{SL}_{2}(k))$ is bad in $G$, while $p$ is good for $G$. For a concrete example, see [15, Rem. 5.12]. This can only happen if $p$ is sufficiently small compared to the rank of $G$, thanks to Theorem 3.3.
(iii). The proofs of Theorems 4.10 and 4.12 by Seitz in [28] depend on separate considerations for each Dynkin type and involve in part intricate arguments for the component groups of centralizers of unipotent elements. In [22], McNinch presents uniform proofs of Seitz’s theorems which are almost entirely free of any case-by-case checks, utilizing methods from geometric invariant theory. However, McNinch’s argument (see [22, Thm. 44]) of the conjugacy result in Theorem 4.10(ii) depends on the fact that for a good $A_{1}$ subgroup $A$ of $G$, the $A$-module $\operatorname{Lie}(G)$ is a tilting module. The latter is established by Seitz in [28, Thm. 1.1].
We present some alternative short and uniform proofs of parts (i)–(iii) of Theorem 4.10 based on the results on associated cocharacters from §2.8. For part (iii) we require the following fact.
Lemma 4.15.
Suppose the connected reductive group $H$ acts on the affine variety $X$. Suppose some Borel subgroup $B$ of $H$ fixes a point $x\in X$. Then $H$ fixes $x$.
Proof.
Because $B$ fixes $x$, there is an induced map $H/B\to X$ given by $hB\mapsto h\cdot x$. Since $H/B$ is projective and irreducible, this map is constant. Hence $H$ fixes $x$. ∎
Proof of Theorem 4.10(i)–(iii).
By the comment before Theorem 4.10, there is no harm in assuming that $G$ is simply connected. According to the Bala–Carter Theorem 2.1, there is a Levi subgroup $L$ of $G$ such that $u$ is distinguished in $L$ (including the case $L=G$). Since $p$ is also good for $L$ (see §2.2), there is an $A_{1}$ subgroup $A$ in $\mathscr{D}L$ containing $u$, owing to Theorem 4.5. Since $p$ is good for $A$, there is a cocharacter $\lambda$ in $\Omega^{a}_{A}(u)$, by Proposition 2.6(i). In particular, $\lambda(\mathbb{G}_{m})$ is a maximal torus in $A$. Then $\lambda$ belongs to $\Omega^{a}_{L}(u)$, by Lemma 2.11 and thanks to Lemma 2.12, it also belongs to $\Omega^{a}_{G}(u)$. It thus follows from Lemma 2.10 that the weights of $\lambda(\mathbb{G}_{m})$ on $\operatorname{Lie}(G)$ do not exceed $22p-$ 2. Thus $A$ is a good $A_{1}$ subgroup of $G$, and so Theorem 4.10(i) follows.
Next let $A,\tilde{A}$ be in $\mathscr{A}(u)$. Let $\lambda\in\Omega^{a}_{A}(u)$ and $\tilde{\lambda}\in\Omega^{a}_{\tilde{A}}(u)$. By the arguments above, we have $\Omega^{a}_{A}(u)\cup\Omega^{a}_{\tilde{A}}(u)\subseteq\Omega^{a}_{G}(u)$. Thanks to Proposition 2.6(ii), $\lambda$ and $\tilde{\lambda}$ are $C_{G}(u)^{\circ}$-conjugate. Further, by Proposition 2.8, we have $C_{G}(u)^{\circ}=C_{G}(\lambda,u)^{\circ}R_{u}(C_{G}(u))$, so that $\lambda$ and $\tilde{\lambda}$ are $R_{u}(C_{G}(u))$-conjugate. There is no harm in assuming that $\lambda=\tilde{\lambda}$. But then $A$ and $\tilde{A}$ share the common Borel subgroup $B=\lambda(\mathbb{G}_{m})U$, where $U$ is the unique $1$-dimensional subgroup in $A$ (and $\tilde{A}$) containing $u$ and normalized by $\lambda(\mathbb{G}_{m})$. It follows from [19, Lem. 2.4] that $A=\tilde{A}$. So Theorem 4.10(ii) holds.
Now let $A$ be in $\mathscr{A}(u)$ and let $\lambda\in\Omega^{a}_{A}(u)$. Then by the argument above, $\lambda\in\Omega^{a}_{G}(u)$. Clearly, $C_{G}(A)\subseteq C_{G}(\lambda,u)$. Let $g\in C_{G}(\lambda,u)$. Then $gAg{{}^{-1}}$ and $A$ share the common Borel subgroup $B=\lambda(\mathbb{G}_{m})U$, as above. Once again, thanks to [19, Lem. 2.4], we conclude that $gAg{{}^{-1}}=A$, and so $g\in N_{G}(A)$. Let $A$ act on $G$ by conjugation. Since $C_{G}(B)=C_{G}(\lambda,u)$ (see (2.7)), $B$ is contained in $C_{A}(g)$. It follows from Lemma 4.15 that $g\in C_{G}(A)$. Consequently, $C_{G}(\lambda,u)\subseteq C_{G}(A)$. Thus we have $C_{G}(\lambda,u)=C_{G}(A)$. Finally, using Proposition 2.8, we get that $C_{G}(u)=C_{G}(A)R_{u}(C_{G}(u))$ is a semidirect product and $C_{G}(A)$ is reductive. Theorem 4.10(iii) follows. ∎
For a short and uniform proof of Theorem 4.10(iv), see [22, Thm. 52].
The following relates the set of cocharacters of $G$ which are associated to a unipotent element $u$ of $G$ of order $p$ to those stemming from good $A_{1}$ overgroups of $u$ in $G$.
Corollary 4.16.
Let $G$ be simple and simply connected. Suppose $p$ is good for $G$ and $u\in G$ is unipotent of order $p$.
  • (i)
    Suppose $u$ is distinguished in the Levi subgroup $L$ of $G$ (see Theorem 2.1). Then any $A_{1}$ subgroup of $L$ containing $u$ belongs to $\mathscr{A}(u)$. In particular, any $A_{1}$ subgroup of $L$ containing $u$ belongs to $\mathscr{A}_{L}(u)$.
  • (ii)
    $\mathscr{A}(u)$ coincides with the set of all $C_{G}(u)^{\circ}$-conjugates of an $A_{1}$ subgroup from part (i).
  • (iii)
    We have a disjoint union
    $\Omega^{a}_{G}(u)=\dot{\bigcup_{A\in\mathscr{A}(u)}}\Omega^{a}_{A}(u).$
  • (iv)
    Let $A\in\mathscr{A}(u)$ and let $\lambda\in\Omega^{a}_{A}(u)$. Then $C_{G}(A)=C_{G}(\lambda,u)$.
  • (v)
    Let $\lambda\in\Omega^{a}_{G}(u)$. Then $C_{G}(\lambda,u)$ is $G$-completely reducible.
  • (vi)
    Let $A\in\mathscr{A}(u)$ and let $\lambda\in\Omega^{a}_{A}(u)$. Then we have
    $C_{G}(u)/C_{G}(u)^{\circ}\cong C_{G}(\lambda,u)/C_{G}(\lambda,u)^{\circ}=C_{G}(A)/C_{G}(A)^{\circ}.$
Proof.
Part (i) is immediate from the first paragraph of the previous proof.
(ii). Let $u$, $L$, and $A\in\mathscr{A}_{L}(u)$ be as in part (i). Then for any $g\in C_{G}(u)^{\circ}$, the conjugate $gAg{{}^{-1}}$ of $A$ is a good $A_{1}$ subgroup of $gLg{{}^{-1}}$ and thus of $G$, as $u$ is also distinguished in $gLg{{}^{-1}}$. Thus any $C_{G}(u)^{\circ}$-conjugate of $A$ is again a good $A_{1}$ subgroup of $G$. Now let $\tilde{A}$ be in $\mathscr{A}(u)$. Theorem 4.10(ii) asserts that there is a $g\in C_{G}(u)^{\circ}$ such that $\tilde{A}=gAg{{}^{-1}}$.
(iii). We first prove that the union in (iii) is disjoint. The argument is more less identical to the proof of Theorem 4.10(ii) above. Let $A,\tilde{A}\in\mathscr{A}(u)$ and let $\lambda\in\Omega^{a}_{A}(u)\cap\Omega^{a}_{\tilde{A}}(u)$. Then $A$ and $\tilde{A}$ share the common Borel subgroup $\lambda(\mathbb{G}_{m})U$, where $U$ is the unique $1$-dimensional subgroup in $A$ and $\tilde{A}$ containing $u$. It follows from [19, Lem. 2.4] that $A=\tilde{A}$.
Thanks to parts (i) and (ii) and the proof of Theorem 4.10(i) above, we get ˙$\dot{\bigcup}_{A\in\mathscr{A}(u)}\Omega^{a}_{A}(u)\subseteq\Omega^{a}_{G}(u)$. Conversely, let $\lambda\in\Omega^{a}_{G}(u)$. By Proposition 2.6(ii), $\lambda$ is $C_{G}(u)^{\circ}$-conjugate to a member of this union.
(iv). This equality is derived in our proof of Theorem 4.10(iii) above.
(v). By part (iii), there is an $A\in\mathscr{A}(u)$ such that $\lambda\in\Omega^{a}_{A}(u)$. Owing to Theorem 4.10(iv), $A$ is $G$-cr, thus so is $C_{G}(A)$, by [6, Cor. 3.17]. The result now follows from part (iv).
(vi). This follows from Proposition 2.8 and (iv). ∎
We note that Corollary 4.16(i) is implicit in [28] and Corollary 4.16(iii) is stated in [22, p. 393].
In [28, §9], Seitz exhibits instances when there is no good $A_{1}$ overgroup of an element of order $p$ when $p$ is bad for $G$. As we explain next, Example 4.2 gives a counterexample to Theorem 4.10(iv) in case $p$ is bad for $G$: that is, it gives a good $A_{1}$ subgroup $A$ containing a unipotent element of order $p$ of $G$ such that $A$ is not $G$-cr. Specifically, we show that some of the $A_{1}$ subgroups in that example are good $A_{1}$ subgroups of $G$, but thanks to Example 4.2, they are not $G$-cr.
Example 4.17 (Example 4.2 continued).
Let $G$ be simple of type $C_{2}$ and $p=2$. Let $\sigma$ be $\operatorname{id}_{G}$ or a $q$-Frobenius endomorphism of $G$. Let $\mathscr{C}$ denote the subregular unipotent class of $G$. Suppose $u\in\mathscr{C}\cap G_{\sigma}$. Then by Example 4.2 there are $\sigma$-stable subgroups $A$ of $G$ of type $A_{1}$ containing $u$ that are not $G$-cr. Specifically, let $E$ be the natural module for $\operatorname{SL}_{2}(k)$. Consider the two conjugacy classes of embeddings of $\operatorname{SL}_{2}(k)$ into $G=\operatorname{Sp}(V)$, where we take either $V\cong E\perp E$ or $V\cong E\otimes E$, as an $\operatorname{SL}_{2}(k)$-module. The images of both embeddings meet the class $\mathscr{C}$ non-trivially. One checks that the highest weight of a maximal torus of $\operatorname{SL}_{2}(k)$ on $\operatorname{Lie}(G)$ is $4$ in the second instance. So in this case the image of $\operatorname{SL}_{2}(k)$ in $G$ is not a good $A_{1}$. In contrast, in the first instance the highest weight of a maximal torus of $\operatorname{SL}_{2}(k)$ on $\operatorname{Lie}(G)$ is $22=2p-$ 2, by Example 4.11. So the image of $\operatorname{SL}_{2}(k)$ in $G$ is a good $A_{1}$ in $\operatorname{SL}(V)$, and so it is a good $A_{1}$ in $G$ as well.
5. Proofs of Theorems 1.1 and 1.31.5
Armed with the results on associated cocharacters from above, we prove Theorems 1.1 and 1.4 simultaneously.
Proof of Theorems 1.1 and 1.4.
First we reduce to the case when $G$ is simple and simply connected. In view of Lemma 3.2, it suffices to show that $H$ (resp. $H_{\sigma}$) is $G$-cr. Thus we need to show that $G$-complete reducibility and the property of being distinguished behave well with respect to the steps in this reduction process. This is achieved by [6, Lem. 2.12] and [15, §4.3], respectively. For, let $\pi\colon G\to G/Z(G)^{\circ}$ be the canonical projection. Owing to [6, Lem. 2.12(ii)(b)] and [15, §4.3], we can replace $G$ with $G/Z(G)^{\circ}$, so without loss we can assume that $G$ is semisimple. Let $G_{1},\ldots,G_{r}$ be the simple factors of $G$. Multiplication gives an isogeny from $G_{1}\times\cdots\times G_{r}$ to $G$. Thus, again by [6, Lem. 2.12(ii)(b)] and [15, §4.3], we can replace $G$ with $G_{1}\times\cdots\times G_{r}$, so we can assume $G$ is the product of its simple factors and moreover that each of these is simply connected. Finally, thanks to [6, Lem. 2.12(i)] and [15, §4.3], it is thus enough to prove the result when $G$ is simple. We may also assume that $H$ is connected and semisimple, since any unipotent element of $H^{\circ}$ is contained in the derived subgroup $\mathscr{D}H^{\circ}$, and $H$ is $G$-ir if $\mathscr{D}H^{\circ}$ is.
Since $u$ has order $p$ also in $H$, there is an overgroup $A$ of $u$ of type $A_{1}$ inside $H$ (resp. $\sigma$-stable $A_{1}$ subgroup of $H$), thanks to Theorem 4.5 in case $p$ is good for $H$, and else by Theorem 4.6 and Lemma 4.8. Since $p$ is good for $A$, there is a cocharacter $\lambda$ in $\Omega^{a}_{A}(u)$, by Proposition 2.6(i). Since $u$ is distinguished in $G$, $\lambda$ belongs to $\Omega^{a}_{G}(u)$, by Lemma 2.11. It follows from Lemma 2.10 that $A$ is a good $A_{1}$ subgroup of $G$ (resp. a good $\sigma$-stable $A_{1}$ subgroup of $G$). Thus, by Theorem 4.10(iv), $A$ is $G$-cr (resp. by Theorem 4.12(iii), $A_{\sigma}$ is $\sigma$-completely reducible and so, thanks to Remark 4.14(i), $A_{\sigma}$ is $G$-cr). Finally, by Lemma 3.2, $A$ (resp. $A_{\sigma}$) is $G$-ir and so is $H$ (resp. $H_{\sigma}$). ∎
As a consequence of Theorems 1.1 and 1.4 we obtain the following.
Corollary 5.1.
Let $G$ be a connected reductive group. Suppose $p$ is good for $G$. Let $\sigma$ be $\operatorname{id}_{G}$ or a $q$-Frobenius endomorphism of $G$. Let $u\in G_{\sigma}$ be unipotent of order $p$. Suppose $u$ is distinguished in the $\sigma$-stable Levi subgroup $L$ of $G$ (see Remark 2.2(ii)). Let $H$ be a $\sigma$-stable connected reductive subgroup of $L$ containing $u$, then $H_{\sigma}$ is $G$-completely reducible.
Proof.
As $p$ is also good for $L$ (see §2.2), it follows from Theorem 1.1 (resp. 1.4) applied to $L$ that $H_{\sigma}$ is $L$-ir and so is $L$-cr. Thus, $H_{\sigma}$ is $G$-cr, by [30, Prop. 3.2]. ∎
Finally, we address Theorems 1.3 and 1.5.
Proof of Theorems 1.3 and 1.5.
By Theorems 1.1 and 1.4, the only cases we need to consider are when $p$ is bad for $G$. If $G$ is classical, then we are in the situation of Lemma 4.1 and Example 4.2.
We are left to consider the case when $G$ is of exceptional type. Then owing to Lemma 4.3, $G$ is of type $G_{2}$ and $p=3$. There is no harm in assuming that $H$ is semisimple. It follows from Example 4.4 that $H$ is $G$-ir. Thus Theorem 1.3 follows. So consider the setting of Theorem 1.5 when $\sigma|_{H}$ is a $q$-Frobenius endomorphism of $H$ in this case. By Corollary 4.7, $u$ belongs to the subregular class of $G_{2}$. It follows from the proof of Lemma 4.3 in [26] that $u$ is contained in a $\sigma$-stable maximal rank subgroup of $G$ of type $A_{1}\tilde{A}_{1}$ and this type is unique. Since $H$ is proper and semisimple, $H\subseteq M$, where $M$ is a $\sigma$-stable maximal rank subgroup of $G$ of type $A_{1}\tilde{A}_{1}$. Since $p$ is good for $H$, there is a $\sigma$-stable subgroup $A$ of $H$ of type $A_{1}$ containing $u$, by Theorem 4.5. Thus $A\subseteq H\subseteq M$. Since $u$ is also distinguished in $M$ and $p=3$ is good for $M$, Theorem 1.4 shows that $A_{\sigma}$ is $M$-ir. Note that $M$ is the centralizer of a semisimple element of $G$ of order $2$ (by Deriziotis’ Criterion, see [10, 2.3]). Since $A_{\sigma}$ is $M$-cr, it is $G$-cr, owing to [6, Cor. 3.21]. Once again, by Lemma 3.2, $A_{\sigma}$ is $G$-ir and so is $H_{\sigma}$. Theorem 1.5 follows. ∎
Remark 5.2.
In [16, §7], Korhonen gives counterexamples to Theorem 1.1 when the order of the distinguished unipotent element of $G$ is greater than $p$ (even when $p$ is good for $G$ [16, Prop. 7.1]). Theorem 3.1 implies that this can only happen when $p<a(G)$. For instances of overgroups of distinguished unipotent elements of $G$ of order greater than $p$ for $p\geq a(G)$ (and $p$ good for $G$), so that Theorem 3.1 applies, see Examples 5.4 and 5.5.
Remark 5.3.
In view of Remark 5.2, it is natural to ask for instances of $G$, $u$ and $H$ when the conclusion of Theorem 3.1 holds even when $p<a(G)$ but $p$ is still good for $G$. If $p$ is good for $G$ and $G$ is simple classical, non-regular distinguished unipotent elements always belong to a maximal rank semisimple subgroup $H$ of $G$, by [36, Prop. 3.1, Prop. 3.2]. For $G$ simple of exceptional type this is also the case in almost all instances of non-regular distinguished unipotent elements, see [36, Lem. 2.1]. Each such $H$ is obviously $G$-irreducible. This is independent of $p$ of course and thus applies in particular when $p<a(G)$. For instance, let $G$ be of type $E_{7}$, $p=5$, and suppose $u$ belongs to the distinguished class $E_{7}(a_{3})$ (resp. $E_{7}(a_{4})$, $E_{7}(a_{5})$). Then $\operatorname{ht}_{J}(\rho)=9$ (resp. $7$, $5$), so $u$ has order $5^{2}$, by Lemma 2.3 in each case. Since $u$ does not have order $5$, Theorem 1.1 does not apply, and since $5<8=a(G)$ neither does Theorem 3.1. Nevertheless, in each case $u$ is contained in a maximal rank subgroup $H$ of type $A_{1}D_{6}$, see [36, p. 52], and each such $H$ is $G$-ir.
We close with several additional higher order examples in good characteristic when Theorem 1.1 does not apply but Theorem 3.1 does.
Example 5.4.
Let $G$ be of type $E_{6}$. Suppose $p$ is good for $G$. In [36, Lem. 2.7], Testerman exhibits the existence of a simple subgroup $H$ of $G$ of type $C_{4}$ whose regular unipotent class belongs to the subregular class $E_{6}(a_{1})$ of $G$. Let $u$ be regular unipotent in $H$. For $p=7$, the order of $u$ is $7^{2}$, by Lemma 2.3, so Theorem 1.1 can’t be invoked to say anything about $H$. However, for $p=7=a(G)$, we infer from Theorem 3.1 that $H$ is $G$-ir.
Example 5.5.
Let $G$ be of type $E_{8}$. Suppose $p=11$. Let $u$ be in the distinguished class $E_{8}(a_{3})$ (resp. $E_{8}(a_{4})$, $E_{8}(b_{4})$, $E_{8}(a_{5})$, or $E_{8}(b_{5})$). From the corresponding weighted Dynkin diagram corresponding to $u$ we get $\operatorname{ht}_{J}(\rho)=17$ (resp. $14$, $13$, $11$, or $11$), see [9, p. 177]. It follows from Lemma 2.3 that in each of these instances $u$ has order $11^{2}$. So we can’t appeal to Theorem 1.1 to deduce anything about reductive overgroups of $u$. But as $911=p\geq a(G)=$ 9, Theorem 3.1 applies and allows us to conclude that each such overgroup is $G$-ir. For example, in each instance, $u$ is contained in a maximal rank subgroup $H$ of $G$ of type $A_{1}E_{7}$ or $D_{8}$, see [36, p. 52].
Acknowledgments: We are grateful to M. Korhonen and D. Testerman for helpful comments on an earlier version of the manuscript. The research of this work was supported in part by the DFG (Grant #RO 1072/22-1 (project number: 498503969) to G. Röhrle). Some of this work was completed during a visit to the Mathematisches Forschungsinstitut Oberwolfach: we thank them for their support. We thank the referee for a number of comments clarifying some points. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.